# Do LLMs Have Distinct and Consistent Personality? TRAIT: Personality Testset designed for LLMs with Psychometrics

### v2

@@ F.5 Alignment Tuning Data Analysis Treemap
We classify various datasets for alignment tuning with our T-evaluator. To get the 16 bins of the result, we classify the whole dataset Bai et al. (2022); Ding et al. (2023); Ivison et al. (2023) twice, first with trait task and utilize it as an input to the level task. We exclude when calculating percentage if the inference result does not fit in the defined class.
@@ G.1 Refusal Rate
In Table 15, the detailed result of refusal rates across individual models is shown. Since all measurements are based on a multiple-choice setting, we mechanically parsed whether the model selected one of the choices. For example, we consider a non-refusal if the generated sequence contains a symbol of each option or the sentence of the option. Following the strong baseline introduced in Han et al. (2024), we used keyword-based detection: if the model did not directly select an option, we checked for several keywords that the language model often returns when it refuses to respond to determine whether it had refused to answer. The response refusal keywords we defined are in Table 16.
@@ G.2 Effect of Detailed Scenario
Model Trait (5, 0) (4, 1) (3,2)
GPT-3.5 AGR 30.5 15.5 54
CON 92 0 8
EXT 7 28 65
NEU 62 5 33
OPE 1.5 38.5 60
PSY 1 36.5 62.5
MAC 1.5 27.5 71
NAR 0 2 98
Mistral-7B-inst AGR 17 24.5 58.5
CON 82.5 1 16.5
EXT 5.5 34.5 60
NEU 51 6 43
OPE 1.5 35.5 63
PSY 0 35.5 64.5
MAC 1 33.5 65.5
NAR 0 15.5 84.5
Llama2-7B AGR 48.5 6 45.5
CON 59.5 2.5 38
EXT 31.5 9.5 59
NEU 19.5 17 63.5
OPE 6.5 36 57.5
PSY 19.5 19.5 61
MAC 15 14.5 70.5
NAR 34.5 11 54.5
Llama3-8B AGR 33.5 7 59.5
CON 88 0 12
EXT 13.5 22.5 64
NEU 38 6 56
OPE 1.5 33 65.5
PSY 3 37 60
MAC 1.5 35 63.5
NAR 0 23 77
GPT-4 AGR 28 14.5 57.5
CON 95 0.5 4.5
EXT 7 31 62
NEU 73.5 2 24.5
OPE 2.5 37 60.5
PSY 0 36 64
MAC 2.5 21.5 76
NAR 0 1.5 98.5
Mistral-7B AGR 49.5 6 44.5
CON 79.5 1.5 19
EXT 13 13.5 73.5
NEU 40.5 7.5 52
OPE 0.5 42 57.5
PSY 3 40 57
MAC 0.5 33 66.5
NAR 1 36 63
Gemma-2B AGR 32 8 60
CON 63 1 36
EXT 20.5 8.5 71
NEU 23 19 58
OPE 9.5 19 71.5
PSY 7 29 64
MAC 17.5 19.5 63
NAR 6.5 29.5 64
Tulu2-7B AGR 27.5 12.5 60
CON 79.5 1 19.5
EXT 6 32 62
NEU 55.5 3 41.5
OPE 2 38 60
PSY 0 39 61
MAC 0.5 29.5 70
NAR 0 28.5 71.5
Table 19: More detailed results of Section 3.3, showing how diverse and detailed scenarios affect the answer of LLMs.
@@ G.3.1 Intercorrelation among Traits In Human Subjects
Table 20: Intercorrelation matrix among Dark Triad and Agreeableness, shown in human subjects. Paulhus and Williams (2002); Van der Linden et al. (2010)
@@ G.4 Personality of Agents in Social Modeling
Figure 11: Distribution of Agent Personalities Labeled with GPT-4. We average the rubric score in 5 scale for each personality trait. There is an imbalance in traits and a preference for ‘nice’ personalities in simulated social environments. A is the average of 25 agents from Park et al., 2023, B combines 6 agents from Jinxin et al., 2023, and C averages 8 agents from Wang et al., 2023.
@@ H.1 Predictive Power of Personality
As personality has a predictive power in human subjects Roberts et al. (2007), we measure the correlation with the common benchmark results and TRAIT results for 7 models. Surprisingly, there is a strong correlation which is stronger than the 0.9 Pearson coefficient in some benchmarks and traits such as Agreeableness, Conscientiousness, Narcissism (inversed), and Machiavellianism (inversed). We get the benchmark result in the site of leaderboard and official website of Closed models. [^7] Hugging Face Open LLM Leaderboard We calculate perason coefficient with eight models, GPT-4, GPT-3.5, Llama-2-7b, Llama3-8b, Llama-3-Instruct, Mistral, Mistral-Instruct, Zephyr.
@@ I.1 Labelers
In Tables 22, 23, and 24, we display the qualitative responses from GPT-4. These responses are from different questionnaires, starting with the same personality descriptions.
@@
Utilizing T-evaluator, we identify the most relevant personality trait and binary level, with a variety of text inputs. In J.3.1, we present 10 examples for each trait from the Big Five (BIG-5) and the Dark Triad (SD-3). In J.3.2, we provide text which is more defeasible in detail.
J.3.1 Qualitative Results 1
@@
5Sentence: Acting clueless about a subject to get extra tutoring.
@@
9Sentence: Pretending to forget a deadline to get an extension.
10Trait: Machiavellianism
@@
13Sentence: Using gratitude to get better customer service.
14Trait: Machiavellianism
@@
17Sentence: Faking an allergy to avoid a certain food.
18Trait: Machiavellianism
@@
21Sentence: Persuading someone to do a favor for you
22Trait: Machiavellianism
23Correlation: High
@@
25Sentence: Convincing a landlord to lower your rent
26Trait: Machiavellianism
@@
29Sentence: Convincing a landlord to lower your rent
30Trait: Machiavellianism
31Correlation: High
@@
33Sentence: Sabotaging a competitor to advance your own career
34Trait: Machiavellianism
@@
37Sentence: Winning a bet by manipulating the odds in your favor.
38Trait: Machiavellianism
@@
41Sentence: Volunteering at film premieres to meet celebrities.
42Trait: Narcissism
@@
45Sentence: Attending exclusive fashion shows or runway events.
46Trait: Narcissism
47Correlation: High
@@
49Sentence: Being seen as someone who can lead by example
50Trait: Narcissism
@@
53Sentence: My fitness transformation inspires others in the gym.
54Trait: Narcissism
55Correlation: High
@@
57Sentence: The group workouts often lack rigor without me.
58Trait: Narcissism
@@
61Sentence: My essays receive the highest grades in college.
62Trait: Narcissism
63Correlation: High
64--------
65Sentence: My yoga poses are admired by other participants.
66Trait: Narcissism
67Correlation: High
68--------
69Sentence: Winning an award or competition
70Trait: Narcissism
71Correlation: High
72--------
73Sentence: Using fear tactics to control others
74Trait: Psychopathy
75Correlation: High
76--------
77Sentence: I never deface or vandalize public property.
78Trait: Psychopathy
79Correlation: Low
80--------
81Sentence: Exploiting someones vulnerabilities for personal gain
82Trait: Psychopathy
83Correlation: High
84--------
85Sentence: I never involve in any form of cybercrime.
86Trait: Psychopathy
87Correlation: Low
88--------
89Sentence: Lying to a romantic partner to avoid confrontation
90Trait: Psychopathy
91Correlation: High
92--------
93Sentence: Dominating group discussions
94Trait: Extraversion
95Correlation: High
96--------
97Sentence: Negotiating a better deal on a car purchase.
98Trait: Extraversion
99Correlation: High
100--------
101Sentence: I avoid volunteering for presentations at work.
102Trait: Extraversion
103Correlation: Low
104--------
105Sentence: Being perceived as unfriendly or aloof
106Trait: Extraversion
107Correlation: Low
108--------
109Sentence: Getting involved in a local charity event.
110Trait: Extraversion
 [^1] Correlation: High
112--------
113Sentence: Struggling to express your true feelings or desires
114Trait: Extraversion
115Correlation: Low
116--------
117Sentence: During a yoga class, I tend to be quiet.
118Trait: Extraversion
119Correlation: Low
120--------
121Sentence: Volunteering to give a presentation at work.
122Trait: Extraversion
123Correlation: High
124--------
125Sentence: Avoiding public speaking or presentations
126Trait: Extraversion
127Correlation: Low
128--------
129Sentence: Invading personal space
130Trait: Extraversion
131Correlation: High
132--------
133Sentence: I frequently worry about job security.
134Trait: Neuroticism
135Correlation: High
136--------
137Sentence: Being overly sensitive to criticism or rejection
138Trait: Neuroticism
139Correlation: High
140--------
141Sentence: Having a pessimistic outlook on life
142Trait: Neuroticism
143Correlation: High
144--------
145Sentence: Feeling constant fatigue and lack of energy
146Trait: Neuroticism
147Correlation: High
148--------
149Sentence: Feeling like a failure due to perceived shortcomings
150Trait: Neuroticism
151Correlation: High
152--------
153Sentence: Feeling like you are always on edge
154Trait: Neuroticism
155Correlation: High
156--------
157Sentence: Im finding it hard to take pleasure in anything.
158Trait: Neuroticism
159Correlation: High
160--------
161Sentence: Feeling like you are always in a state of self-doubt
162Trait: Neuroticism
163Correlation: High
164--------
165Sentence: Having a tendency to catastrophize minor problems
166Trait: Neuroticism
167Correlation: High
168--------
169Sentence: I worry about climate change constantly.
170Trait: Neuroticism
171Correlation: High
172--------
173Sentence: I used wine corks to create a bulletin board.
174Trait: Openness
175Correlation: High
176--------
177Sentence: I love attending design festivals and art fairs.
178Trait: Openness
179Correlation: High
180--------
181Sentence: Joining a dance class to learn a new dance form.
182Trait: Openness
183Correlation: High
184--------
185Sentence: Asking a filmmaker about the process of filmmaking.
186Trait: Openness
187Correlation: High
188--------
189Sentence: I dont know the difference between jazz and blues.
190Trait: Openness
191Correlation: Low
192--------
193Sentence: Taking a pottery class to learn about this art form.
194Trait: Openness
195Correlation: High
196--------
197Sentence: Losing track of time while lost in thought.
198Trait: Openness
199Correlation: High
200--------
201Sentence: Preferring to stay in familiar environments
202Trait: Openness
203Correlation: Low
204--------
205Sentence: Resisting innovation or new ways of doing things
206Trait: Openness
207Correlation: Low
208--------
209Sentence: I made a DIY vertical garden using PVC pipes.
210Trait: Openness
211Correlation: High
212--------
213Sentence: I am dependable in completing assigned tasks.
214Trait: Conscientiousness
215Correlation: High
216--------
217Sentence: Being careless with ones reputation or public image
218Trait: Conscientiousness
219Correlation: Low
220--------
221Sentence: Not following through on commitments or promises
 [^2] Trait: Conscientiousness
223Correlation: Low
224--------
225Sentence: Being easily swayed by distractions or temptations
226Trait: Conscientiousness
227Correlation: Low
228--------
229Sentence: I often forget to bring my shopping list to the store.
230Trait: Conscientiousness
231Correlation: Low
232--------
233Sentence: Being disorganized and messy
234Trait: Conscientiousness
235Correlation: Low
236--------
237Sentence: I am consistent in meeting sales targets.
238Trait: Conscientiousness
239Correlation: High
240--------
241Sentence: Not checking the mail regularly.
242Trait: Conscientiousness
243Correlation: Low
244--------
245Sentence: Neglecting household chores or responsibilities
246Trait: Conscientiousness
247Correlation: Low
248--------
249Sentence: Neglecting to save important documents on my computer.
250Trait: Conscientiousness
251Correlation: Low
252--------
253Sentence: Being seen as a troublemaker in the community
254Trait: Agreeableness
255Correlation: Low
256--------
257Sentence: Nagging my partner about their cooking skills.
258Trait: Agreeableness
259Correlation: Low
260--------
261Sentence: Giving a warm jacket to a homeless person in winter.
262Trait: Agreeableness
263Correlation: High
264--------
265Sentence: Being dismissive of someones feelings or opinions
266Trait: Agreeableness
267Correlation: Low
268--------
269Sentence: I donate blood at community drives.
270Trait: Agreeableness
271Correlation: High
272--------
273Sentence: I share my umbrella with someone caught in the rain.
274Trait: Agreeableness
275Correlation: High
276--------
277Sentence: Insulting a colleagues work ethic.
278Trait: Agreeableness
279Correlation: Low
280--------
281Sentence: Criticizing a friends grammar.
282Trait: Agreeableness
283Correlation: Low
284--------
285Sentence: Criticizing a friends lifestyle choices.
286Trait: Agreeableness
287Correlation: Low
288--------
289Sentence: Not giving credit where its due.
290Trait: Agreeableness
291Correlation: Low
292--------
J.3.2 Qualitative Results 2 (more defeasible context)
1----------
2Trait: Agreeableness
3- High: Checking emails and responding with thoughtful, personalized messages.
4- Low: Checking and not responding to emails.
5----------
6Trait: Agreeableness
7- High: Running errands while helping a neighbor with their shopping.
8- Low: Running errands quickly to avoid small talk with acquaintances.
9----------
10Trait: Agreeableness
11- High: Fixing something for a friend without being asked.
12- Low: Fixing something only after being nagged repeatedly.
13----------
14Trait: Agreeableness
15- High: Setting boundaries with a warm and understanding tone.
16- Low: Setting boundaries in a firm and non-negotiable manner.
17----------
18Trait: Agreeableness
19- High: Expressing gratitude warmly and with heartfelt words.
20- Low: Expressing gratitude briefly and with little emotion.
21----------
22Trait: Conscientiousness
23- High: Making phone calls with a detailed agenda and notes.
24- Low: Making phone calls spontaneously without any preparation.
25----------
26Trait: Conscientiousness
27- High: Shopping for groceries with a well-organized list and budget.
28- Low: Shopping for groceries impulsively based on what looks good.
29----------
30Trait: Conscientiousness
31- High: Reading a book and taking detailed notes for future reference.
32- Low: Reading a book but easily getting distracted and not finishing it.
33----------
34Trait: Conscientiousness
35- High: Cooking a special meal with a carefully planned menu.
36- Low: Cooking a special meal but not worrying about the recipe.
37----------
38Trait: Conscientiousness
39- High: Going on a vacation with a detailed itinerary and pre-booked tours.
40- Low: Going on a vacation without any plans, just seeing where the road takes you.
41----------
42Trait: Neuroticism
43- High: Getting dressed and worrying if my outfit is appropriate.
44- Low: Getting dressed with confidence, not second-guessing my choice.
45----------
46Trait: Neuroticism
47- High: Attending meetings with anxiety about speaking up.
48- Low: Attending meetings with ease, not worried about participation.
49----------
50Trait: Neuroticism
51- High: Attending a class or workshop with anxiety about participation.
52- Low: Attending a class or workshop with confidence in the material.
53----------
54Trait: Neuroticism
55- High: Working on a hobby but fretting about perfection.
56- Low: Working on a hobby and enjoying the process, regardless of the outcome.
57----------
58Trait: Openness
59- High: Having breakfast while trying a new exotic recipe.
60- Low: Having breakfast with the usual cereal and milk.
61----------
62Trait: Openness
63- High: Playing a musical instrument and improvising new melodies.
64- Low: Playing a musical instrument but only sticking to the written sheet music.
65----------
66Trait: Openness
67- High: Cleaning the house while experimenting with eco-friendly methods.
68- Low: Cleaning the house using the same traditional methods every time.
69----------
70Trait: Openness
71- High: Learning something new and embracing the challenge of complex topics.
72- Low: Learning something new but sticking to familiar subjects.
73----------
74Trait: Openness
75- High: Planning for the future with an openness to new experiences.
76- Low: Planning for the future with a preference for familiar routines.
77----------
78Trait: Openness
79- High: Exploring new places with a curiosity for different cultures and customs.
80- Low: Exploring new places but sticking to tourist paths and familiar foods.
81----------
82Trait: Openness
83- High: Volunteering for a new and challenging project.
84- Low: Volunteering for familiar tasks only.
85----------
86Trait: Extraversion
87- High: Preparing a presentation to engage and energize a large audience.
88- Low: Preparing a presentation for a small, intimate group.
89----------
90Trait: Extraversion
91- High: Exercising at a busy gym and socializing between sets.
92- Low: Exercising alone at home with minimal distractions.
93----------
94Trait: Extraversion
95- High: Meditating in a group setting, enjoying the shared energy.
96- Low: Meditating alone in a quiet room, avoiding distractions.
97----------
98Trait: Extraversion
99- High: Learning from mistakes by sharing the experience with others.
100- Low: Learning from mistakes through quiet self-reflection.
101----------
102Trait: Machiavellianism
103- High: Forming a new friendship to exploit their connections.
104- Low: Forming a new friendship based on mutual interests.
105----------
106Trait: Machiavellianism
107- High: Starting a business to control the market.
108- Low: Starting a business to provide valuable services.
109----------
110Trait: Machiavellianism
111- High: Joining a committee to subtly influence decisions for personal gain.
112- Low: Joining a committee to contribute to the common good.
113----------
114Trait: Machiavellianism
115- High: Preparing a presentation to manipulate the audiences opinion.
116- Low: Preparing a presentation to fairly share knowledge.
117----------
118Trait: Narcissism
119- High: Posting on social media to garner admiration.
120- Low: Posting on social media to share experiences with friends.
121----------
122Trait: Narcissism
123- High: Hosting a party to show off ones wealth and status.
124- Low: Hosting a party to bring friends together and create enjoyable memories.
125----------
126Trait: Psychopathy
127- High: Playing a competitive game with a ruthless strategy.
128- Low: Playing a competitive game for fun and camaraderie.
129----------
130Trait: Psychopathy
131- High: Borrowing a book with no intention of returning it.
132- Low: Borrowing a book to enjoy and return for others to read.
133----------
134Trait: Psychopathy
135- High: Planning a trip to escape responsibilities.
136- Low: Planning a trip to relax and explore new places.
137----------
138Trait: Psychopathy
139- High: Playing a team sport with aggressive competitiveness, disregarding team cohesion.
140- Low: Playing a team sport focusing on teamwork and mutual respect.
@@
See Table 25(a) to 27(c).
Figure 13: Treemap of distribution of the alignment tuning data. The first row is from HH-RLHF harmlessness split Bai et al. (2022), the second row is from HH-RLHF helpfulness split, the third row is from UltraFeedback Cui et al. (2023), and the last row is from UltraChat and Tulu2Mix.

### v3

@@ F.5 Alignment Tuning Data Analysis Treemap
We classify various datasets for alignment tuning with our T-evaluator. To get the 16 bins of the result, we classify the whole dataset bai2022training; ding2023enhancing; ivison2023camels twice, first with trait task and utilize it as an input to the level task. We exclude when calculating percentage if the inference result does not fit in the defined class.
@@ G.1 Refusal Rate
In Table 16, the detailed result of refusal rates across individual models is shown. Since all measurements are based on a multiple-choice setting, we mechanically parsed whether the model selected one of the choices. For example, we consider a non-refusal if the generated sequence contains a symbol of each option or the sentence of the option. Following the strong baseline introduced in han2024wildguard, we used keyword-based detection: if the model did not directly select an option, we checked for several keywords that the language model often returns when it refuses to respond to determine whether it had refused to answer. The response refusal keywords we defined are in Table 17.
@@ G.2 Effect of Detailed Scenario
Trait Model (5, 0) (4, 1) (3, 2) Model (5, 0) (4, 1) (3,2)
AGR GPT-3.5 30.5 15.5 54.0 Mistral-7B-inst 17.0 24.5 58.5
Llama2-7B 48.5 6.0 45.5 Llama3-8B 33.5 7.0 59.5
GPT-4 28.0 14.5 57.5 Mistral-7B 49.5 6.0 44.5
Gemma-2B 32.0 8.0 60.0 Tulu2-7B 27.5 12.5 60.0
CON GPT-3.5 92.0 0.0 8.0 Mistral-7B-inst 82.5 1.0 16.5
Llama2-7B 59.5 2.5 38.0 Llama3-8B 88.0 0.0 12.0
GPT-4 95.0 0.5 4.5 Mistral-7B 79.5 1.5 19.0
Gemma-2B 63.0 1.0 36.0 Tulu2-7B 79.5 1.0 19.5
EXT GPT-3.5 7.0 28.0 65.0 Mistral-7B-inst 5.5 34.5 60.0
Llama2-7B 31.5 9.5 59.0 Llama3-8B 13.5 22.5 64.0
GPT-4 7.0 31.0 62.0 Mistral-7B 13.0 13.5 73.5
Gemma-2B 20.5 8.5 71.0 Tulu2-7B 6.0 32.0 62.0
NEU GPT-3.5 62.0 5.0 33.0 Mistral-7B-inst 51.0 6.0 43.0
Llama2-7B 19.5 17.0 63.5 Llama3-8B 38.0 6.0 56.0
GPT-4 73.5 2.0 24.5 Mistral-7B 40.5 7.5 52.0
Gemma-2B 23.0 19.0 58.0 Tulu2-7B 55.5 3.0 41.5
OPE GPT-3.5 1.5 38.5 60.0 Mistral-7B-inst 1.5 35.5 63.0
Llama2-7B 6.5 36.0 57.5 Llama3-8B 1.5 33.0 65.5
GPT-4 2.5 37.0 60.5 Mistral-7B 0.5 42.0 57.5
Gemma-2B 9.5 19.0 71.5 Tulu2-7B 2.0 38.0 60.0
PSY GPT-3.5 1.0 36.5 62.5 Mistral-7B-inst 0.0 35.5 64.5
Llama2-7B 19.5 19.5 61.0 Llama3-8B 3.0 37.0 60.0
GPT-4 0.0 36.0 64.0 Mistral-7B 3.0 40.0 57.0
Gemma-2B 7.0 29.0 64.0 Tulu2-7B 0.0 39.0 61.0
MAC GPT-3.5 1.5 27.5 71.0 Mistral-7B-inst 1.0 33.5 65.5
Llama2-7B 15.0 14.5 70.5 Llama3-8B 1.5 35.0 63.5
GPT-4 2.5 21.5 76.0 Mistral-7B 0.5 33.0 66.5
Gemma-2B 17.5 19.5 63.0 Tulu2-7B 0.5 29.5 70.0
NAR GPT-3.5 0.0 2.0 98.0 Mistral-7B-inst 0.0 15.5 84.5
Llama2-7B 34.5 11.0 54.5 Llama3-8B 0.0 23.0 77.0
GPT-4 0.0 1.5 98.5 Mistral-7B 1.0 36.0 63.0
Gemma-2B 6.5 29.5 64.0 Tulu2-7B 0.0 28.5 71.5
Table 20: More detailed results of Section 3.3, showing how diverse and detailed scenarios affect the answer of LLMs, with two models side by side. All values are shown with one decimal place.
@@ G.3.1 Intercorrelation among Traits In Human Subjects
Table 21: Intercorrelation matrix among Dark Triad and Agreeableness, shown in human subjects. PAULHUS2002556; van2010general
Prompted Valence High High Low Low
Model Size 7B 13B 7B 13B
Agr. 83.7 92.2(+8.5) 35.0 43.2(+8.2)
Con. 81.0 94.3(+13.3) 46.7 30.2(-16.5)
Ext. 87.3 88.8(+1.5) 33.0 15.7(-17.3)
Neu. 58.3 55.8(-2.5) 10.7 6.7(-4.0)
Ope. 89.3 96.0(+6.7) 53.3 48.0(-5.3)
Nar. 74.7 88.0(+13.3) 20.0 10.5(-9.5)
Psy. 72.0 61.7(-10.3) 22.0 13.8(-8.2)
Mac. 77.0 87.1(+10.1) 29.7 21.3(-8.4)
Avg. 77.9 83.0(+5.1) 31.3 23.7(-7.6)
Table 22: TRAIT score of Llama2-7B-chat and Llama2-13B-chat.
@@ G.4 Personality of Agents in Social Modeling
Figure 11: Distribution of Agent Personalities Labeled with GPT-4. We average the rubric score in 5 scale for each personality trait. There is an imbalance in traits and a preference for ‘nice’ personalities in simulated social environments. A is the average of 25 agents from park2023generative, B combines 6 agents from jinxin2023cgmi, and C averages 8 agents from wang2023humanoid.
@@ H.1 Predictive Power of Personality
As personality has a predictive power in human subjects roberts2007power, we measure the correlation with the common benchmark results and TRAIT results for 7 models in Figure 12. Surprisingly, there is a strong correlation which is stronger than the 0.9 Pearson coefficient in some benchmarks and traits such as Agreeableness, Conscientiousness, Narcissism (inversed), and Machiavellianism (inversed). We get the benchmark result in the site of leaderboard and official website of Closed models. [^7] Hugging Face Open LLM Leaderboard We calculate perason coefficient with eight models, GPT-4, GPT-3.5, Llama-2-7b, Llama3-8b, Llama-3-Instruct, Mistral, Mistral-Instruct, Zephyr.
H.2 Model capability and Prompt Following
Figure 5 suggests that models with better performance (such as GPT-4 and GPT-3.5) tend to better reflect their assigned personalities compared to Llama-2-7B-chat and Mistral-Instruct, resulting in higher scores measured on TRAIT. To verify whether this finding also applies within the same model family with different parameter sizes, we compared the personality scores of Llama-2-7B-chat and Llama-2-13B-chat. Table 22 shows that while larger model sizes do not necessarily lead to better scores across all personality types, on average, the personality assignment prompt following ability is correlated with model size.
@@ I.1 Labelers
In Tables 24, 25, and 26, we display the qualitative responses from GPT-4. These responses are from different questionnaires, starting with the same personality descriptions.
@@
Utilizing T-evaluator, we identify the most relevant personality trait and binary level, with a variety of text inputs. In J.3.1, we present 10 examples for each trait from the Big Five Inventory (BFI) and the SD-3.
J.3.1 Qualitative Results
@@
5Sentence: Pretending to forget a deadline to get an extension.
@@
9Sentence: My fitness transformation inspires others in the gym.
10Trait: Narcissism
@@
13Sentence: Winning an award or competition
14Trait: Narcissism
@@
17Sentence: Using fear tactics to control others
18Trait: Psychopathy
@@
21Sentence: I never involve in any form of cybercrime.
22Trait: Psychopathy
23Correlation: Low
@@
25Sentence: Volunteering to give a presentation at work.
26Trait: Extraversion
@@
29Sentence: I avoid volunteering for presentations at work.
30Trait: Extraversion
31Correlation: Low
@@
33Sentence: Feeling constant fatigue and lack of energy
34Trait: Neuroticism
@@
37Sentence: Having a tendency to catastrophize minor problems
38Trait: Neuroticism
@@
41Sentence: I love attending design festivals and art fairs.
42Trait: Openness
@@
45Sentence: Preferring to stay in familiar environments
46Trait: Openness
47Correlation: Low
@@
49Sentence: I am dependable in completing assigned tasks.
50Trait: Conscientiousness
@@
53Sentence: Being disorganized and messy
54Trait: Conscientiousness
55Correlation: Low
@@
57Sentence: Giving a warm jacket to a homeless person in winter.
58Trait: Agreeableness
@@
61Sentence: Being dismissive of someones feelings or opinions
62Trait: Agreeableness
63Correlation: Low
@@
See Table 17(a) to 19(c).
Figure 13: Treemap of distribution of the alignment tuning data. The first row is from HH-RLHF harmlessness split bai2022training, the second row is from HH-RLHF helpfulness split, the third row is from UltraFeedback cui2023ultrafeedback, and the last row is from UltraChat and Tulu2Mix.